what are the steps to find the equation of the indicated line . and writing i in slope intercept form .
y-intercept(0,3) slope 1

Answer by mananth(16946)

m= 1

Plug value of the slope and point ( 0,3 ) in
Y = m x + b
3.00 = 0 + b
b= 3 - 0
b= 3
So the equation will be
Y =x + 3
m.ananth@hotmail.ca